# Prediction-Powered Adaptive Shrinkage Estimation

### v1

@@ 7 Conclusion
This paper introduces PAS, a novel method for compound mean estimation that effectively combines PPI and EB principles. We motivate the problem through the lens of variance reduction and contextual prior information—then demonstrate how PAS achieves both goals, in theory and in practice. Our paper differs from many other PPI-related works in its focus on estimation, so a natural next step is to develop average coverage controlling intervals for the means centered around PAS. To this end, it may be fruitful to build on the robust empirical Bayes confidence intervals of Armstrong et al. (2022). Modern scientific inquiries increasingly demand the simultaneous analysis of multiple related problems. The framework developed in this paper—combining empirical Bayes principles with machine learning predictions—represents a promising direction for such settings.
Appendix A The Correlation-Aware Unbiased Risk Estimate
Theorem A.1.
@@ Appendix A The Correlation-Aware Unbiased Risk Estimate
Remark A.2 (Connection to SURE).
Stein’s Unbiased Risk Estimate (SURE) was proposed in Charles Stein’s seminal work [1981] to study the quadratic risk in Gaussian sequence models. As a simple special case of SURE, let $Z\sim\mathcal{N}(\theta,\,\sigma^{2})$ and let $h:\mathbb{R}\to\mathbb{R}$ be an absolutely continuous function and $\mathbb{E}_{\theta}\!\left[|h^{\prime}(Z)|\right]<\infty$, then SURE is defined as
@@
with the property that $\mathbb{E}_{\theta}\!\left[\text{SURE}(h)\right]=R(h(Z),\theta)=\mathbb{E}_{\theta}[(h(Z)-\theta)^{2}]$. A proof of this argument relies on Stein’s lemma, an identity specific to Gaussian random variables Stein [1981]. Now consider the specific linear shrinkage estimator $h_{c}(Z):=cZ+(1-c)Y$, with $c\in[0,1]$ and $Y\in\mathbb{R}$ being fixed (that is, $Y$ is a constant, or $Y$ is independent of $Z$ and we condition on $Y$). Then $\mathrm{SURE}$ takes the following form:
@@
where in ()($\star$)( ⋆ ) we used the fact that in this case (with $Y$ fixed or $Y$ independent of $Z$), it holds that $\gamma=0$ so that the definition of $\mathrm{CURE}$ in (19) simplifies. This explains how CURE defined in A.1 is connected to SURE.
We make one last remark: The derivation of SURE itself requires Gaussianity. However, for linear shrinkage rules as $h_{c}(Z)$, SURE only depends on the first two moments of the distribution of $Z$ and thus is an unbiased estimator of quadratic risk under substantial generality as long as $\mathbb{E}_{\theta}[Z]=\theta$ and $\mathrm{Var}_{\theta}[Z]=\sigma^{2}$. This remark has been made by previous authors, e.g., Kou and Yang [2017], Ignatiadis and Wager [2019] and is important for the assumption-lean validity of PAS.
Appendix B Proofs of Theoretical Results
B.1 Proof of Theorem. 4.1
For each problem $j\in[m]$, we are shrinking the PT estimator $\hat{\theta}^{\mathrm{PT}}_{j}$ obtained from the first stage towards $\tilde{Z}^{f}_{j}$, the prediction mean on the unlabeled data. Conditioning on $\eta_{j}$, we denote
where all the first and second moments of $\hat{\theta}_{j}^{\mathrm{PT}}$ and $\tilde{Z}^{f}_{j}$ exist under the conditions of 2.2. For each global $\omega\geq 0$, the shrinkage parameter for the $j$-th problem is defined as $\omega_{j}:=\omega/(\omega+\tilde{\sigma}^{2}_{j})$. Then, following the result in Theorem A.1, CURE for $\hat{\theta}_{j,\omega_{j}}^{\mathrm{PAS}}:=\omega_{j}\hat{\theta}_{j}^{\mathrm{PT}}+(1-\omega_{j})\tilde{Z}^{f}_{j}$,
is an unbiased estimator of the risk, i.e.,
B.2 Formal conditions and proof of Proposition 5.1

### v2

@@ 8 Conclusion
This paper introduces PAS, a novel method for compound mean estimation that effectively combines PPI and EB principles. We motivate the problem through the lens of variance reduction and contextual prior information—then demonstrate how PAS achieves both goals, in theory and in practice. Our paper differs from many other PPI-related works in its focus on estimation, so a natural next step is to develop average coverage controlling intervals for the means centered around PAS. To this end, it may be fruitful to build on the robust empirical Bayes confidence intervals of Armstrong et al. (2022). Modern scientific inquiries increasingly demand the simultaneous analysis of multiple related problems. The framework developed in this paper represents a promising direction for such settings.
Acknowledgments
We thank Dan Kluger and Qingqi Zhang for thoughtful feedback on an earlier version of our manuscript. We also thank Claire Donnat, Yichen Ji, Lihua Lei, Aaron Schein, and Tijana Zrnic for helpful discussions. We are grateful to Bennett Hunter and the compute cluster at the Data Science Institute for supporting our computational resources. We thank all four reviewers for their constructive feedback. N.I. gratefully acknowledges support from NSF (DMS 2443410).
Impact Statement
By developing a method to improve the efficiency and accuracy of mean estimation using machine learning predictions, this research has the potential to enhance data analysis across various domains where labeled data is limited but predictive models are available.
We acknowledge that advancements in machine learning can have broader societal consequences. However, the ethical considerations directly arising from this methodological contribution are those typically associated with the general advancement of statistical methods and machine learning. We do not foresee specific negative ethical impacts unique to this work that require further detailed discussion.
Appendix A Further Connections and Intuitions Behind PAS
A.1 Further Connections to Related Works
In this part, we elaborate on two important connections to existing work.
Stratified PPI (StratPPI; Fisch et al. (2024)).
Fisch et al. (2024) also consider a setting with stratification of observations into subproblems. However, their overall aim is still to estimate a single parameter, rather than multiple parameters, as we do. Thus, in solving individual problems, for instance, they still adhere to the requirement of unbiasedness. As an example, let us revisit the Galaxy Zoo 2 application. Therein, we mentioned the following distinction between our work and previous methods in the PPI literature:
  • Estimate the fraction $\theta$ of spiral galaxies across the whole universe (Angelopoulos et al., 2023).
  • (Our work) Estimate the fraction $\theta_{j}$ of spiral galaxies within the $j$-th cluster of galaxies for $j\in[m]$.
Now suppose that galaxies in the whole universe can be partitioned into the $m$ clusters above. Then the goal of StratPPI is:
Estimate the fraction $\theta$ of spiral galaxies across the whole universe by proceeding with an intermediate step that involves estimating $\theta_{1},\dotsc,\theta_{m}$.
We continue to make this connection explicit in our mean estimation setting. Suppose (dropping subscripts for convenience) that we start with covariate-outcome pairs $(X,Y)\in\mathcal{X}\times\mathcal{Y}$ distributed as
$(X,Y)\,{\buildrel\text{iid}\over{\sim}\,}\mathbb{P}.$
Moreover suppose that there exist pairwise disjoint strata $\mathcal{A}_{1},\dotsc,\mathcal{A}_{m}$ whose union is the full covariate space $\mathcal{X}$, that is, $\mathcal{X}=\dot{\bigcup}_{j=1}^{m}\mathcal{A}_{j}$. Then define,
$w_{j}:=\mathbb{P}[X\in\mathcal{A}_{j}],\;\;\quad\theta_{j}:=\mathbb{P}[Y\mid X\in\mathcal{A}_{j}],\;\;\quad j\in[m],$
and observe the key equality for $\theta:=\mathbb{E}[Y]$,
$\theta=\sum_{j=1}^{m}w_{j}\theta_{j}.$
StratPPI assumes that the probabilities $w_{j}$ are known exactly and then samples $n_{j}$ labeled as well as $N_{j}$ unlabeled observations from the conditional distribution $\mathbb{P}[\cdot\mid X\in\mathcal{A}_{j}]$. [^15] Part of the contribution of StratPPI includes a strategy for allocating resources and choosing $n_{j}$ and $N_{j}$ for different problems with a view toward minimizing the variance of the final estimator of $\theta$. These samples are then used alongside PPI++ to estimate $\theta_{j}$ by $\hat{\theta}^{\mathrm{PT}}_{j}$, almost verbatim to the approach we described in Section 4.1. Finally, StratPPI estimates $\theta$ via:
$\hat{\theta}^{\mathrm{SPPI}}:=\sum_{j=1}^{m}w_{j}\hat{\theta}^{\mathrm{PT}}_{j}.$
To summarize, the settings of StratPPI and our paper are similar, but the goals and methods differ. StratPPI seeks an unbiased estimate of a single parameter $\theta$, while we seek to estimate a parameter vector $(\theta_{1},\dotsc,\theta_{m})$ as well as possible in mean squared error, while allowing for the possibility of bias. Moreover, the asymptotic regimes in these two papers are different: StratPPI keeps $m$ fixed and takes $n_{j},N_{j}\to\infty$, while we assume that the number of problems grows $(m\to\infty)$ while $n_{j},N_{j}$ are allowed to remain bounded.
PPI++ (Angelopoulos et al., 2024) with multivariate mean estimand.
Angelopoulos et al. (2024) (and other papers in the PPI literature) consider statistical problems going beyond the estimation of a univariate mean. For instance, the results of Angelopoulos et al. (2024) also accommodate multivariate mean estimands. Here we explain how one can frame our compound mean estimation setting into the estimation of a single but multivariate mean estimand, which then fits in the more generalized formulation for power tuning considered in PPI++.
If we assume $n_{j}=n$ and $N_{j}=N$ for all $j$, we can collate the responses and predictions from all problems into $(f(X_{i.}),Y_{i.})_{i=1}^{n}$ and $(f(\tilde{X}_{i.}))_{i=1}^{N}$, where
$Y_{i.}=(Y_{i1},...,Y_{im})^{\intercal},\quad f(X_{i.})=(f(X_{i1}),...,f(X_{im}))^{\intercal},\quad f(\tilde{X}_{i.})=(f(\tilde{X}_{i1}),...,f(\tilde{X}_{im}))^{\intercal}.$
By independence across problems, we have the covariance structures $\mathrm{Var}[f(\tilde{X}_{i.})]=\text{diag}((\tau_{j}^{2})_{j=1}^{m})$ and $\mathrm{Cov}[f(X_{i.}),Y_{i.}]=\text{diag}((\gamma_{j})_{j=1}^{m})$ where $\text{diag}(\cdot)$ constructs a diagonal matrix with its arguments on the diagonal. Angelopoulos et al. (2024) then consider a class of estimators indexed by a single weighting parameter $\lambda$ that is applied to the entire vector:
$\hat{\theta}^{\text{PPI}}_{\lambda}:=\frac{1}{n}\sum_{i=1}^{n}Y_{i.}+\lambda\left(\frac{1}{N}\sum_{i=1}^{N}f(\tilde{X}_{i.})-\frac{1}{n}\sum_{i=1}^{n}f(X_{i.})\right).$
The optimal $\lambda^{*}$ is chosen by minimizing the trace of an asymptotic covariance $\Sigma^{\lambda}$. For the mean estimation problem, $\Sigma^{\lambda}$ simplifies to the covariance of $\hat{\theta}^{\text{PPI}}_{\lambda}$ and can be exactly calculated in finite-sample setting (without relying on $n,N\to\infty$). Minimizing $\Sigma^{\lambda}$ over $\lambda$ thus admits a simple closed-from solution:
$\lambda^{*}:=\frac{n^{-1}\text{Tr}(\mathrm{Cov}[f(X_{i.}),Y_{i.}])}{\frac{n+N}{nN}\text{Tr}(\mathrm{Var}[f(\tilde{X}_{i.})])}=\frac{n^{-1}\sum_{j=1}^{m}\gamma_{j}}{\frac{n+N}{nN}\sum_{j=1}^{m}\tau_{j}^{2}}$ (19)
Going back to our original formulation with multiple problems, what we are doing here is to perform power tuning across problems to minimize the sum of MSEs—which equals the trace of the covariance matrix $\text{Tr}(\mathrm{Var}[f(\tilde{X}_{i.})])$ in the PPI++ formulation—over a single (univariate) $\lambda$. Finally, if $n_{j}$ and $N_{j}$ vary across $j$, the above PPI++ formulation fails to hold, but the idea to power tune all problems together carries over to our multiple problem setting. The univariate optimal tuning parameter takes the same form as Equation 19:
$\lambda^{*}:=\frac{\sum_{j=1}^{m}n_{j}^{-1}\gamma_{j}}{\sum_{j=1}^{m}\frac{n_{j}+N_{j}}{n_{j}N_{j}}\tau_{j}^{2}}.$
In fact, this choice of $\lambda^{*}$ leads to a new estimator, denoted as UniPT, that will be further explored in Appendix C.
A.2 Further Intuitions Behind the Design of PAS
A.2.1 Why Does Sharing Information Across Problems Help?
Here we provide an explanation of the fundamental statistical difference between estimating a single mean versus estimating a lot of means ($m\to\infty$) with an eye toward providing intuition of how we can share information across problems. We emphasize that this is only meant for intuition; the empirical Bayes literature, starting with James & Stein (1961), provides a more nuanced picture.
Imagine a simplified setting with only a single problem ($m=1$, as in the standard PPI setting). We further assume that $N_{j}=\infty$, so that $\tilde{Z}^{f}_{1}\equiv\mathbb{E}_{\eta_{1}}[f(\tilde{X}_{1})]$ and we also assume that $\tilde{\sigma}_{1}^{2}$ is known. Moreover, suppose we want to choose between the PT estimator $\hat{\theta}_{1}^{\text{PT}}$ and the prediction mean $\tilde{Z}_{1}^{f}$ by comparing their MSEs in a data-driven way. We know that the PT estimator is unbiased and its MSE is $\tilde{\sigma}^{2}_{1}$, but we cannot estimate $\mathbb{E}[(\theta_{1}-\tilde{Z}_{1}^{f})^{2}]$ accurately since we only have a single $\theta_{1}$ (a single problem). At best, we can use $(\hat{\theta}_{1}^{\text{PT}}-\tilde{Z}_{1}^{f})^{2}-\tilde{\sigma}_{1}^{2}$ as an unbiased estimate of this quantity since $\mathbb{E}_{\eta_{1}}[(\hat{\theta}_{1}^{\text{PT}}-\tilde{Z}_{1}^{f})^{2}]-\tilde{\sigma}_{1}^{2}=\mathbb{E}_{\eta_{1}}[(\theta_{1}-\tilde{Z}_{1}^{f})^{2}].$
Now suppose we have multiple problems, then we can more precisely learn how good the prediction mean $\tilde{Z}_{j}^{f}$ is for estimating $\theta_{j}$ by repeating the above estimation procedure for each problem and averaging the results. This average turns out to be not only unbiased but also consistent. To wit, as $m\to\infty$, by the law of large numbers,
$\frac{1}{m}\sum_{j=1}^{m}\left((\hat{\theta}_{j}^{\text{PT}}-\tilde{Z}_{j}^{f})^{2}-\tilde{\sigma}^{2}_{j}\right)\stackrel{{\scriptstyle\mathbb{P}}}{{\to}}\mathbb{E}_{\mathbb{P}_{\eta}}[(\theta_{j}-\tilde{Z}_{j}^{f})^{2}],$ (20)
where we emphasize that the mean squared on the right-hand side also integrates with respect to the meta-distribution $\smash{\mathbb{P}_{\eta}}$ that determines the distribution of the $\smash{\theta_{j}}$ (cf. Assumption 2.1). This illustrates a mechanism for sharing information across problems: in a first step, we can consistently estimate the average squared difference between the true parameters $\theta_{j}$ and the prediction means $\tilde{Z}_{j}^{f}$ (i.e., how good the ML predictor is on average) by aggregating information from many parallel problems. Then, based on the first step, we can choose whether to use $\boldsymbol{\tilde{Z}}^{f}$ or $\boldsymbol{\hat{\theta}}^{\mathrm{PT}}$ according to their MSE.
Our actual implementation is more involved but embodies the same fundamental principle: we construct a consistent estimate for the MSE of a family of shrinkage estimators. While other factors such as heterogeneity in sample sizes ($n_{j},N_{j}$) introduce further complexities, the core idea of sharing information here carries over to the overall design and justification of PAS.
A.2.2 Analogy between $Z_{j}^{f}$ and $\mathbb{E}[\theta\mid\psi]$
In Equation 16, we informally draw an analogy between the prediction mean $Z_{j}^{f}$ in the compound PPI problem and the posterior mean $\mathbb{E}[\theta\mid\psi]$ in the stylized Gaussian model. Our goal here is to provide a heuristic motivation for the one-dimensional parameterized family of weights $\omega_{j}$ in Equation 15. We emphasize that the ultimate success of this parameterization choice is judged by the empirical results.
Our heuristic motivation building on the stylized example of Section 3 is as follows. We seek to estimate $\theta$ by taking a convex combination $\hat{\theta}^{cl}-\xi$ and a shrinkage target $s$,
$w(\hat{\theta}^{cl}-\xi)\,+\,(1-w)s,$
for $w\in[0,1]$. Our main point is that for several choices of shrinkage target $s$, the best weight $w$ can be written in the form
$w=\frac{\omega}{\omega+\sigma_{\varepsilon}^{2}},$ (21)
[…]
  • The Bayes estimator with contextual prior in Equation 12 uses $s=\mathbb{E}[\theta\mid\phi]$ and weight $w$ as in (21) with
    $\omega=\mathbb{E}[(\theta-\mathbb{E}[\theta\mid\phi])^{2}]=\mathbb{E}[\mathrm{Var}[\theta\mid\phi]]=\sigma_{\theta\mid\phi}^{2}.$
  • The Bayes estimator without contextual prior uses $s=\mathbb{E}[\theta]=0$ and weight $w$ as in (21) with
    $\omega=\mathbb{E}[(\theta-\mathbb{E}[\theta])^{2}]=\mathrm{Var}(\theta)=\sigma_{\theta}^{2}.$
  • Suppose now that we ask for the best convex combination (not necessarily a Bayes predictor) between $\hat{\theta}^{cl}-\xi$ and $s=h(\phi)$ where $h(\cdot)$ is some fixed function. Then, the weight $w$ minimizing MSE can be shown to take the form in (21) with
    $\omega=\mathbb{E}[(\theta-h(\phi))^{2}]=\mathbb{E}[\mathrm{Var}[\theta\mid\phi]]+\mathbb{E}[(h(\phi)-\mathbb{E}[\theta\mid\phi])^{2}].$
    The above expression is interesting as it forces us to inflate “$\omega$”, i.e., to shrink less toward $h(\phi)$ in a way that depends on how close $h(\phi)$ is to $\mathbb{E}[\theta\mid\phi]$. See Ignatiadis & Wager (2019, last paragraph of Section 3) and Section A.2.3 below for further discussion of this point.
The takeaway is that for lots of possible predictors, the optimal weights have the same parameterized form up to the single parameter $\omega$ that varies according to the quality of the predictor. This motivates our one-dimensional family of weights. Once this family has been motivated, we learn $\omega$ in Equation 15 in a way that does not depend on the above analogy at all by minimizing CURE.
A.2.3 Elaboration on Proposition 5.3
Suppose (as in Proposition 5.3) that $n_{j}$ is the same across all problems, $N_{j}=\infty$, and that second moments $\rho_{j},\tau_{j},\sigma_{j}$ are identical across all problems. Then we could ask: what is the best convex combination between $\hat{\theta}_{j}^{\text{PT}}$ and $\tilde{Z}_{j}^{f}$ in the following sense:
$\omega_{j}^{*}\in\operatorname*{arg\,min}_{\omega_{j}\geq 0}\mathbb{E}_{\mathbb{P}_{\eta}}\big{[}\{\theta_{j}-(\omega_{j}\hat{\theta}_{j}^{\text{PT}}+(1-\omega_{j})\tilde{Z}_{j}^{f})\}^{2}\big{]}.$
By direct calculation (note that the right-hand side is a convex quadratic in $\omega_{j}$) we find that:
$\omega_{j}^{*}=\frac{\mathbb{E}_{\mathbb{P}_{\eta}}[(\theta_{j}-\tilde{Z}_{j}^{f})^{2}]}{\mathbb{E}_{\mathbb{P}_{\eta}}[(\theta_{j}-\tilde{Z}_{j}^{f})^{2}]+\tilde{\sigma}^{2}}.$
This implies the following intuitive result: the larger the MSE $\mathbb{E}_{\mathbb{P}_{\eta}}[(\theta_{j}-\tilde{Z}_{j}^{f})^{2}]$, the less weight we should assign to $\tilde{Z}_{j}^{f}$. If we evaluate the MSE at this optimal $\omega_{j}^{*}$, we recover precisely the upper bound of Proposition 5.3.
Appendix B The Correlation-Aware Unbiased Risk Estimate
Theorem B.1.
@@ Appendix B The Correlation-Aware Unbiased Risk Estimate
Remark B.2 (Connection to SURE).
Stein’s Unbiased Risk Estimate (SURE) was proposed in Charles Stein’s seminal work (1981) to study the quadratic risk in Gaussian sequence models. As a simple special case of SURE, let $Z\sim\mathcal{N}(\theta,\,\sigma^{2})$ and let $h:\mathbb{R}\to\mathbb{R}$ be an absolutely continuous function and $\mathbb{E}_{\theta}\!\left[|h^{\prime}(Z)|\right]<\infty$, then SURE is defined as
@@
with the property that $\mathbb{E}_{\theta}\!\left[\text{SURE}(h)\right]=R(h(Z),\theta)=\mathbb{E}_{\theta}[(h(Z)-\theta)^{2}]$. A proof of this argument relies on Stein’s lemma, an identity specific to Gaussian random variables (Stein, 1981). Now consider the specific linear shrinkage estimator $h_{c}(Z):=cZ+(1-c)Y$, with $c\in[0,1]$ and $Y\in\mathbb{R}$ being fixed (that is, $Y$ is a constant, or $Y$ is independent of $Z$ and we condition on $Y$). Then $\mathrm{SURE}$ takes the following form:
@@
where in ()($\star$)( ⋆ ) we used the fact that in this case (with $Y$ fixed or $Y$ independent of $Z$), it holds that $\gamma=0$ so that the definition of $\mathrm{CURE}$ in (22) simplifies. This explains how CURE defined in B.1 is connected to SURE.
We make one last remark: The derivation of SURE itself requires Gaussianity. However, for linear shrinkage rules as $h_{c}(Z)$, SURE only depends on the first two moments of the distribution of $Z$ and thus is an unbiased estimator of quadratic risk under substantial generality as long as $\mathbb{E}_{\theta}[Z]=\theta$ and $\mathrm{Var}_{\theta}[Z]=\sigma^{2}$. This remark has been made by previous authors, e.g., Kou & Yang (2017); Ignatiadis & Wager (2019) and is important for the assumption-lean validity of PAS.
Appendix C Details on PAS with Unknown Second Moments
In this appendix we explain how to apply PAS when second moments are unknown. In Section C.1 we describe sample-based estimators of the second moments. We also develop UniPT (Section C.2) and UniPAS (Section C.3), two new estimators that extend our framework to scenarios where second moments are unknown and must be estimated from data. We derive these methods and their theoretical guarantees within the PAS asymptotic regime, where the number of problems $m\to\infty$ while individual sample sizes $n_{j},N_{j}$ remain bounded. Specifically, UniPT introduces a global power-tuning strategy, and UniPAS builds upon it to perform adaptive shrinkage, both without requiring knowledge of true second-moment parameters.
Notations.